OpenStudy (anonymous):

(cos Θ + cos Θ)2 + (cos Θ + cos Θ)2 Can someone please help me with this??

OpenStudy (astrophysics):

Are you sure it's \[(\cos \theta+\cos \theta)^2+(\cos \theta+\cos \theta)^2\]

OpenStudy (anonymous):

yeah you have it right

OpenStudy (astrophysics):

Ok, so notice that \[\cos \theta + \cos \theta = 2\cos \theta\]

OpenStudy (anonymous):

okay

OpenStudy (astrophysics):

\[(2\cos(\theta))^2+(2\cos(\theta))^2\] so we have this now

OpenStudy (astrophysics):

What do you get

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

sin^2theta?

OpenStudy (astrophysics):

No just follow algebra rules what is \[(2\cos \theta)^2\]

OpenStudy (anonymous):

(2cosT)(2cost)?

OpenStudy (astrophysics):

Yeah

OpenStudy (astrophysics):

So what is that

OpenStudy (anonymous):

cos^2t?

OpenStudy (astrophysics):

\[(2\cos(\theta))^2 = 4\cos^2(\theta)\]

OpenStudy (anonymous):

ohhh okay

OpenStudy (astrophysics):

Now you have \[4\cos^2 \theta+4\cos^2 \theta\] what is that

OpenStudy (astrophysics):

It's just like saying 4x+4x

OpenStudy (anonymous):

8cos^2t?

OpenStudy (astrophysics):

Yup

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and then what

OpenStudy (astrophysics):

That's it

OpenStudy (anonymous):

thats not one of the options

OpenStudy (anonymous):

unless i can just put 8?

OpenStudy (astrophysics):

What are the options

OpenStudy (anonymous):

sin2 Θ 4 8 cos2 Θ

OpenStudy (astrophysics):

When you post a question, give everything

OpenStudy (astrophysics):

I think you gave the wrong problem, because all of those are wrong.

OpenStudy (anonymous):

okay im sorry

OpenStudy (anonymous):

(cos Θ + cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (anonymous):

thats the problem

OpenStudy (astrophysics):

Can you take an image of it

OpenStudy (astrophysics):

All your options are incorrect :\

OpenStudy (anonymous):

those are the options i was given

OpenStudy (astrophysics):

It should be 8 cos^2x

OpenStudy (astrophysics):

theta*

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

those are the answers i was given im sorry

OpenStudy (astrophysics):

Is there more to the problem

OpenStudy (anonymous):

no

OpenStudy (astrophysics):

Well, I'm not sure then, who ever made your problem must've been lazy.

OpenStudy (anonymous):

i thik so thanks though

OpenStudy (astrophysics):

Np

OpenStudy (astrophysics):

You should message your teacher